# Career Modeling with Missing Data and Traces

### v2

@@ 3 Public-Private paths in French elite civil servants
In this second part of the article, we showcase the general methodology on real data taken from the social and political sciences: public-private paths of civil servants in France. While our focus is on the inferential process itself, we also discuss the applied literature and extend it in several manners.
3.1 Presentation of the studies
3.1.1 General framing
Paths of civil servants in the private sector in France have been studied by multiple authors (Bouzidi et al., 2010; Kolopp, 2021; Rouban, 2014) and have been a source of interest in the general public. We explore two aspects assessed as interesting in the literature but not yet described, or not described with such precision: how the structure of public-private paths has evolved since 1990 (Bouzidi et al., 2010); and how patterns of these paths vary across many organizations (Charle, 1987) or individual characteristics like gender (Kolopp, 2021). To this end, we will study two populations of elite civil servants: former students of the National School of Administration (École Nationale d’Administration, hereafter ENA) in Section 3.3, and members of powerful services (listed below) in Section 3.4. Previous studies have focused on detailed data sources only: well described career paths for a small number of people. As such, they have not explored thoroughly these questions, which require large amounts of data, and very diverse data.
We review the relevant applied literature in sections 3.3 and 3.4. We first introduce our data sources along some particularities of the analysis. Details for how we constructed and assembled the population can be found in Appendix A.
3.1.2 Detailed and trace data sources
Recall that we combine two data sources: a detailed data set for a smaller number of individuals, and a trace data set which covers the entire population but with less detail. Quantitative details on the size of each population and their coverage in the detailed source are provided in sections 3.3.1, 3.4.1 and in Appendix B.
Our detailed data come from the website LinkedIn. LinkedIn is a social media platform oriented towards professional networking, where users describe their past work experience on their profile. Many managers, both in the public and private sectors, have a profile, but not the entirety of the population. In addition, as these data are auto-biographical, individuals may choose to not include all relevant information: there are gaps in profiles, which means that there might be missing data even for individuals with a LinkedIn profile. This means that even when an individual is in this detailed data source, it can still be useful to complement the analysis with a trace data source.
Our trace data source is the Official Journal of the French Republic (hereafter JO or “Journal Officiel”), an administrative record with a significant legal value that includes new pieces of regulation, executive decisions and nominations. This source documents nominations for many top-level positions in public organizations. It also includes, in certain cases, information about promotions, other career changes (such as retirement) and decisions taken by authorities that might involve named civil servants. Importantly, certain key positions are systematically described, and this source can thus be used to define exhaustive and representative populations. However, not all public positions are described, and there is often no duration attached to these positions, which explains why it may only be treated as a traces dataset. The source was digitized starting in 1990, and this defines our temporal frame (1990-2022).
3.1.3 Specificities in the implementation
@@
Indices and span of the analysis. We consider the trajectories of individuals through time $t=1,...,T$, where an increment by one unit corresponds to 6 months: $t=1$ corresponds to January-June 1990, and $T=68$ corresponds to July-December 2022. Each individual $i$ is modeled on an interval $[T_{min}^{i},T_{\max}^{i}]$. We let $T_{\min}^{i}$ be their first membership-defining trace (see Appendix A). If the individual has a trace that states they retired, we use that date for $T_{\max}^{i}$; else we let $T_{\max}^{i}=T$, where $T$ corresponds to our last observation time (July-December 2022). Individuals are thus modeled on different spans, a fact that is possible in the general framework but makes notation more cumbersome.
State spaces and parameters. As our aim is to describe public-private paths, we study the state space $\mathcal{X}=\{0,1\}$, with 00 corresponding to the public sector and $1$ to the private sector. It is possible to describe heterogeneity of the public sector by including information in the covariates W and without increasing the size of the state space. We thus have two parameters of interest indexed by individuals and time: $\gamma_{i,t}^{0}:=\gamma_{i,t}^{0\rightarrow 1}$ the probability to exit the public sector at time $t$, and $\gamma_{i,t}^{1}:=\gamma_{i,t}^{1\rightarrow 0}$ the probability for an individual who had previously left to the private sector to return to the public sector. Finally, we describe traces using simply a dichotomous variable as this led to better results: we take $\mathcal{Y}=\{0,1\}$, with $Y_{i,t}=1$ if we observe at least one trace in JO in the given six-months interval and $Y_{i,t}=0$ otherwise. This is modeled as $Y_{i,t}\sim\text{Bernoulli}(\lambda_{i,t})$.
@@
Configuration for autoregressive traces. We code three variables to describe the autoregressive aspect of traces. Two are given by $A_{i,t}(\varphi)=({\sum_{0<s\leq t}\varphi^{s}Y_{t-s}^{i}})/({\sum_{0<s\leq t}\varphi^{s}})$, where we set $\varphi=1$ to prioritize long-term information and $\varphi=0.8$ for medium-term information. $A_{i,t}(1)$ and $A_{i,t}(0.8)$ take values between 00 and $1$. For short-term information, we create a transformation of the time lapsed since the last trace
@@
Handling of boundary effects. An issue is that when $t$ is low, there are not many past trace variables. This becomes problematic because a specificity of our setting is that we use traces both to define the population (as the starting positions we are interested in are systematically described) and as a complementary source (for all traces). This means that by construction, $Y_{i,T^{i}_{\min}}=1$, which needs to be taken into account. As a solution, we ignore this first trace (setting the value to zero), and include specific covariates to handle points before any traces were found: $\textbf{1}_{\{A_{t,i}(1)=0\}}$, and $L_{t,i}\times\textbf{1}_{\{A_{t,i}(1)=0\}}$. A more complex handling is required when including time-sensitive coefficients, and we include instead an interaction between B-splines for age and the indicator $\textbf{1}_{\{A_{t,i}(1)=0\}}$ (see below, models 3, 4, 5 of Study 1).
@@
Prior distributions and convergence. Results presented throughout are with non-informative, flat (improper) prior distributions. Other tests were performed with low-informative priors, e.g. $\mathcal{N}(0,\sigma^{2})$ with $\sigma=5$ for model 3 in study 1, with unchanged results. Contrary to fully unsupervised methods where the dependency on the prior choice could be higher, our setting only exhibits a small dependency, as points taken from the detailed data source are not simulated an thus reduce the relative importance played by the prior in the posterior distribution. We run our MCMC sampler on 5000 iterations and discard the first 200 as a burn-in period. Visual checks show that this is sufficient to guarantee convergence and the MCMC sampler mixes well. Throughout studies 1 and 2, the ESS of all parameters is above 200, and the $\hat{R}$ below $1.006$, indicating that the Markov chain has converged and has appropriately explored the posterior distribution.
3.2 Empirical demonstrations on synthetic data
We present empirical demonstrations of multiple claims made in Section 2, taking advantage of the notation and intuitions that arise from this specific case. We aim to demonstrate two claims: in a reasonable case with a well-specified model, researchers can expect to gain substantial statistical power by adopting our procedure; and when using a standard HMM despite having heterogeneity in the trace-emitting process, errors there communicate and lead to erroneous conclusions for interest parameters, further motivating our approach.
Gains of statistical power. For the first claim, we simulate trajectories using the following form for parameters, with $g^{-1}$ the inverse logit transform:
$\forall i,t,\ \gamma^{0\rightarrow 1}_{i,t}=g^{-1}(-4+C_{i}),\ \gamma^{1\rightarrow 0}_{i,t}=g^{-1}(-4),\ \lambda_{i,t}=g^{-1}(-1.5),$ (6)
with $C_{i}\sim\mathcal{U}([0,1])$ an individual observed variable, that could correspond in our example to the cohort of individual $i$. Our main parameter of interest is the impact of the cohort on $\gamma_{0}$. These coefficients are similar to estimates provided for the simplest models in Section 3.3, so this simulation experiment can be deemed realistic. Each trajectory is initiated at state 00, and is of length 60. We simulated three datasets, with $N=200,500,$ and $1000$ individuals, where everyone is observed in the first two cases (and so no data augmentation is performed), but only 200 individuals in the third case.
Table 1: Reporting on the first synthetic data example
Case 1 Case 2 Case 3 True value
n 200 500 1000
$n_{\text{obs}}$ 200 500 200
$\gamma^{0}$, Intercept [-4.23,-3.52] [-4.25,-3.81] [-4.2,-3.81] -4
$\gamma^{0}$, Cohort [-1.05,1.25] [0.16,1.61] [0.74,1.98] 1
$\gamma^{1}$ [-4.92,-4.05] [-4.1,-3.69] [-4.34,-3.84] -4
$\lambda$ [-1.54,-1.41] [-1.52,-1.44] [-1.5,-1.44] -1.5
We report on the concentration of the posterior of coefficients, by reporting on 90% credibility intervals in Table LABEL:synth:ex1, from which different conclusions can be drawn. First, case 3 performs better than case 1, especially for the cohort coefficient for $\gamma^{0}$, where the posterior is much more concentrated on the true value. Second, cases 2 and 3 can be said to have relatively similar performances, with comparable posteriors and uncertainty. However, the gain here is that instead of observing 500 individuals, only 200 individuals had to be observed in the detailed data source, and imprecise knowledge using traces compensated precise knowledge on a population more than twice larger. The exact population size numbers will vary from application to application, the main idea being that the more “informative” the traces are, the more useful the data augmentation procedure will turn out to be.
Issues of standard HMM misspecification. For the second claim, we simulate trajectories using the following form for parameters
$\forall i,t,\ \gamma^{0\rightarrow 1}_{i,t}=g^{-1}\left(-4+B_{i,t}\right),\ \gamma^{1\rightarrow 0}_{i,t}=g^{-1}(-4),\ \lambda_{i,t}=g^{-1}(-2+2.5A_{i,t}(0.8)),$ (7)
where $B_{i,t}=\frac{t}{60}$ is a time-varying covariate, which here could correspond to age or period, and $A_{i,t}(0.8)$ the autocorrelation covariate as defined earlier. We simulate 2000 sequences of length 60 and observe 30% of them. We implement two models: one well-specified (ours), and one misspecified (standard HMM), where covariate $A_{i,t}$ is not included.
Table 2: Reporting on the second synthetic data example
Well-specified Misspecified True value
$\gamma^{0}$, Intercept [-4.28,-3.98] [-3.74,-3.49] -4
$\gamma^{0}$, Age [0.86,2.02] [-0.3,0.7] 1
$\gamma^{1}$, Intercept [-4.35,-3.95] [-4.1,-3.79] -4
$\lambda$, Intercept [-2.03,-1.98] [-1.49,-1.45] -2
$\lambda$, $A(0.8)$ [2.4,2.6] 2.5
Coefficients are reported in Table LABEL:synth:ex2. In the well-specified case, the output credibility intervals correctly contain the true parameter values. On the other hand, in the misspecified HMM case, the coefficient for the age effect $B_{i,t}$ is wrong in sign and magnitude. Such problems will arise for time-varying covariates when an autocorrelation behavior is not taken into account in the modeling phase. The same point is also shown by the cross-validation reconstruction losses obtained in Study 1 below.
3.3 Study 1. ENA graduates and temporary exit
3.3.1 Motivation and population definition
The National School of Administration ENA was founded in 1945 to select and train high civil servants for the French state. Graduates of this school (around 80 to 100 a year) play a central role in the political and administrative life of the country, and they constitute a natural population when studying elite civil servants (Bouzidi et al., 2010; Rouban, 2014). We study students admitted between 1990 and 2019, before the school was replaced by the National Institute for Public Service (INSP) in 2021111Data taken from Journal Officiel were first assembled and cleaned up by Nathann Cohen and put up on an API on the website steinertriples. Throughout the article, we rely on the amazing meta-data provided by this API, in addition to our own data cleaning procedures.. We only keep track of individuals who graduated from ENA and whose first position was working for the state, which we operationalize by only keeping individuals for which at least one trace can be found in the three years following their graduation. The number of students identified using this method by year of admission can be found in Appendix B and forms an exhaustive total of $N=2715$.
ENA graduates constitute a natural population when studying public-private career paths, as they hold a central position in the organization of the French state. Rouban (2014) conducted a survey over 6 student year-groups ($N=620$) who graduated in 1969-1970, 1989-90, and 1999-2000. He reports that, 10 years after graduating, 8.3% of students from the first two cohorts were working in a private company, compared to 14.5% for the two next cohorts, and 9.5% for the last two, which indicates year-to-year variation but not a clear trend towards increasing exit rate to the private sector.
The results of Rouban (2014) seem to show that the average time taken for an ENA graduate before leaving the public sector has drastically decreased over time, which would imply a stark increase in the exit rate from the state. We must however note that their methodology does not allow for comparing different cohorts. Indeed, many statistics considered in this study, like average time to leave to the private sector, are subject to right-censoring; this was not accounted for or acknowledged. This is problematic as the observation frames are significantly different across cohorts: it was, at the time of publication, 44 years for the 1970 cohort, but only 14 years for the 2000 cohort. We assume the difference with our results is due to our methodology properly taking into account the different lengths of observation frames. We conclude that the variation reported by Rouban (2014) is artefactual. The other quantitative conclusions of Rouban (2014) are not subject to this bias, and are coherent with our results.
Bouzidi et al. (2010) focus on a subgroup of ENA graduates, of the “civil administrator” status working in the Finance Ministry, between 1960 and 2002 ($N=438$), using hand-collected data. They show that, on their sample, about 40% of individuals went at least once into the private sector 20 years after graduating. The advantage of their approach is that they used a formal duration analysis model and are not susceptible to right-censoring biases. They however do not attempt to estimate whether there was an acceleration or not on their period of study, and stress that doing such inference is difficult for later cohorts because of right-censoring. They also estimate time spent in the private sector using simulations using two Weilbull distributions, one for the public sector and one for the private sector. Our work can be seen as a discretization of such a procedure that also enables missing data imputation and inclusion of covariates.
@@ 3.3.1 Motivation and population definition
We build our first study from these surveys and provide two improvements. First, thanks to our methodology, we can broaden the scope and analyze all students who graduated between 1991 and 2021, including the 55% do not have a LinkedIn profile, as well as the individuals who have a partially incomplete LinkedIn profile. In addition, since our modeling accounts allows individuals to leave for the private sector and then return to the public sector, we can measure more precisely the public exit rate, and distinguish between individuals the public sector permanently or temporarily.
@@ 3.3.2 Models
We implement four initial configurations which are described here from simplest to most complex. Models 1 and 2 are introduced for pedagogical purposes. Model 1 is a standard hidden Markov model: the covariate vector $W_{i,t}$ only includes the intercept. Model 2 is a Markov switching model for which only the autocorrelation covariates described in section 3.1.3 are included for the estimation of $\lambda$.
Models 3, 4 (and 5) include time-specific interest covariates and notably age. Age is encoded as time elapsed since the individual was admitted to ENA, on a range from 0 to 1 (with 1 corresponding to the entire time frame, 32 years). Age is modeled in a semi-parametric fashion, using B-splines of degree 2, with 5 degrees of freedom including an intercept, and boundary knots in 0 and 1, leading to $g=1,...,5$ covariates $S_{i,t}^{g}$ in the linear model (with the R package splines). We do this in order to saturate the model on the age dimension, so we can safely identify an effect related to cohort or period effects. Using splines works better than multiple powers of age: under this parameterization, parameters are less correlated, so the sampler mixes faster. As the spline basis already includes an intercept, no general intercept is included. Recall from our discussion in Section 3.1.3 that we have to use a more complex scheme to handle boundary effects for $\lambda$, and interact these five covariates with $\mathbb{I}_{A(1)=0}$ for all three models. More specifically, Model 3 adds a time-varying component depending on the age of individuals for $\gamma_{0}$ and $\gamma_{1}$. Model 4 adds a cohort effect for the estimation of $\gamma^{0}$, $\gamma^{1}$, and $\lambda$. Cohort is encoded as the time at which the individual was admitted at ENA, ranging between 0 and 1.
Finally, we showcase that it is possible to answer questions that are less descriptive and more centered on substantive social mechanisms by studying the dependence between public-private paths and electoral cycles. We implement an additional Model 5, which is similar to Model 4, except for the fact that instead of a cohort effect, we estimate a period effect using a dummy variable $P_{i,t}$ with value $1$ if $t$ or $t+1$ is an election time point, and 00 else. Election time points correspond to Jan-Jun 1995, 2002, 2007, 2012, 2017, 2022 (French presidential elections). This dummy is shifted by one unit (so $t+1$ or $t+2$) for $\lambda$.
@@ 3.3.3 Analysis of Model Accuracy
We first analyse the output for models 1, 2, and 3. We study both the posterior of parameters for simpler models and the quality of data augmentation by performing analyses where detailed data source information is left out for some observed individuals.
Parameters posterior. The simplicity of models 1 and 2 enables full reporting on the posterior of their coefficients, which we do in Table 3 with their 0.05, 0.50 and 0.95 quantiles. Full information about model coefficients can be found in Appendix D.
model $M_{1}$ $M_{2}$
@@
$\gamma_{1}$, (Intercept) -4.054 -3.938 -3.83 -3.667 -3.525 -3.391
$\lambda$, (Intercept) -1.51 -1.492 -1.474 -1.853 -1.786 -1.718
$\lambda$, $L$0.16 -0.134 -0.108
$\lambda$, $A(1)$0.133 0.088 0.315
@@
$\lambda$, $I(A(1)=0):L$0.173 -0.137 -0.101
Table 3: Coefficients for Models 1 and 2.
Comparison between models 1 and 2 shows why there is a need for autoregressive coefficients and to handle boundary effects, as discussed in section 3.1.3. The main effect of adding autocorrelation coefficients from Model 1 to Model 2 can be observed in the difference between their median intercept for $\lambda$. For Model 1, the median intercept is at $-1.492$ ($I_{90}=[-1.51,-1.474]$), which essentially corresponds to an average probability of 0.22 for an individual to generate a trace at any point in time, which is higher than the one obtained for Model 2, at $-1.786$ ($I_{90}=[-1.853,-1.718]$), and which is more in line with what could be expected. All credible regions for autocorrelation coefficients except $A(1)$ do not contain zero despite that they are quite correlated by construction. Finally, the coefficient responsible for handling the boundary effect $\mathbb{I}_{A(1)=0}$ is very different from zero and negative, showing indeed that it is needed. Configurations without it also have very different values for all other coefficients.
Evaluation by reconstruction loss. We repeat the inference under models 1, 2, 3 but with some data withheld. In this experiment, we exclude from the data 20% of individuals in the detailed data source. We can thus assess the data augmentation quality could on points $(i,t)\in S_{\text{obs.,exc.}}$ that are observed but excluded from the analysis. We adopt a simple cross-entropy loss on the trajectory for such points $(i,t)\in S_{\text{obs.,exc.}}$, $l_{i,t}=-(X_{i,t}^{\text{obs}}\log(X_{i,t}^{\text{pred}})+(1-X_{i,t}^{\text{obs}})\log(1-X_{i,t}^{\text{pred}})$), where predictions are computed as the average of observations generated in the Gibbs step over the last $300$ iterations of the chain, with probability clipping at 0.01 and 0.99, and observations are taken from the detailed data source. For Model 1, we also implement a version where no information about the traces was used at all, which we denote Model 1b. Full reporting can be found in Appendix D.
The average cross-entropy loss obtained for the different models is 0.39 for Model 1, 0.29 for Model 1b, 0.22 for Model 2, and 0.21 for Model 3. A first very clear conclusion is that HMMs without autoregressive components are not suited for the task, as Model 1b – which did not even have access to traces! – performed better data reconstruction than Model 1. We observe however that better reconstruction is obtained once including autoregressive components (in models 2 and 3). The change in accuracy from Model 2 to Model 3 is small in size, but our interest is rather to perform hypothesis testing when including the additional age coefficients. Overall, this indicates that one could obtain better reconstruction by adopting more complex models, but this would be at the cost of model interpretability and hypothesis testing, which is our main focus.

### v3

@@ 5 Public-Private paths in French elite civil servants
We now return to our problem of interest and apply this method to public-private paths of French civil servants. We first focus on the population of ENA graduates in Section 5.2, then consider all top civil servants in Section 5.3.
5.1 Specificities in the implementation
@@
Indices and span of the analysis. We consider the trajectories of individuals through time t=1,,Tt=1,...,T, where an increment by one unit corresponds to 6 months: $t=1$ corresponds to January-June 1990, and $T=68$ corresponds to July-December 2022. Each individual $i$ is modeled on an interval $T_{min}^{i},T_{\max}^{i}]$. We let $T_{\min}^{i}$ be their first membership-defining trace (see Appendix A), which is typically the beginning of their career in the civil service. If the individual has a trace that states they retired, we use that date for $T_{\max}^{i}$; else we let $T_{\max}^{i}=T$, where $T$ corresponds to our last observation time (July-December 2022). Individuals are thus modeled on different spans, a fact that is possible in the general framework but makes notation more cumbersome.
@@
Configuration for autoregressive traces. We code three variables to describe the autoregressive aspect of traces. Two are given by $A_{i,t}(\varphi)=({\sum_{0<s\leq t}\varphi^{s}Y_{t-s}^{i}})/({\sum_{0<s\leq t}\varphi^{s}})$, where we set $\varphi=1$ to prioritize long-term information and $\varphi=0.8$ for medium-term information. $A_{i,t}(1)$ and $A_{i,t}(0.8)$ take values between 0 and $1$. For short-term information, we create a transformation of the time lapsed since the last trace
@@
Handling of boundary effects. An issue is that when $t$ is low, there are not many past trace variables. This becomes problematic because a specificity of our setting is that we use traces both to define the population (as the starting positions we are interested in are systematically described) and as a complementary source (for all traces). This means that by construction, $Y_{i,T^{i}_{\min}}=1$, which needs to be taken into account. As a solution, we ignore this first trace (setting the value to zero), and include specific covariates to handle points before any traces were found: $\textbf{1}_{\{A_{t,i}(1)=0\}}$, and $L_{t,i}\times\textbf{1}_{\{A_{t,i}(1)=0\}}$. A more complex handling is required when including time-sensitive coefficients, and we include instead an interaction between B-splines for age and the indicator $\textbf{1}_{\{A_{t,i}(1)=0\}}$ (see below, models 3, 4, 5 in Section 5.2.1).
@@
Prior distributions and convergence. Results presented throughout are with non-informative, flat (improper) prior distributions. Other tests were performed with low-informative priors, e.g. $\mathcal{N}(0,5^{2})$ for model 3 in Section 5.2.1, with unchanged results. Contrary to fully unsupervised methods where the dependency on the prior choice could be higher, our setting only exhibits a small dependency, as points taken from the detailed data source are not simulated and thus reduce the relative importance played by the prior in the posterior distribution. We run our Metropolis-within-Gibbs sampler for 5 0005\,000 Gibbs iterations with $M_{\text{Metropolis}}=30$ steps inside each iteration; we discard the first 200 as a burn-in period. Visual checks show that this is sufficient to guarantee convergence and the MCMC sampler mixes well. For all analyses, the ESS of all parameters is above 200, and the $\hat{R}$ is below $1.006$, indicating that the Markov chain has converged and has appropriately explored the posterior distribution. We repeated some analyses with 50 00050\,000 Gibbs iterations; these longer runs yielded similar results, confirming that the MCMC samplers have run for enough iterations.
5.2 Subpopulation analysis: ENA graduates and temporary exit
@@
We draw inspiration from the surveys of rouban_norme_2014 and bouzidi_pantouflage_2010 and first focus on civil servants who graduated from ENA. Our methodology provides two improvements. First, thanks to our methodology, we can broaden the scope and analyse all students who graduated within our time frame, including the 45% do not have a LinkedIn profile, as well as the individuals who have a partially incomplete LinkedIn profile. In addition, since our modeling allows individuals to leave for the private sector and then return to the public sector, we can measure more precisely the public exit probability, and distinguish between individuals leaving the public sector permanently or temporarily.
We consider students admitted between 1990 and 2019, before the school was replaced by the National Institute for Public Service (INSP) in 2021. We only keep track of individuals who graduated from ENA and whose first position was working for the state, which we operationalize by only keeping individuals for which at least one trace can be found in the three years following their graduation. The number of students identified using this method by year of admission can be found in Appendix B and forms an exhaustive total of $N=2715$. To our knowledge, the most substantive existing analysis of ENA graduate paths is that of rouban_norme_2014, who focused on 6 year groups ($N=620$): our framework allows us to study a substantially larger population.
@@ 5.2.1 Models
We implement four initial configurations which are described here from simplest to most complex. Models 1 and 2 are introduced for pedagogical purposes. Model 1 is a standard hidden Markov model: the covariate vector $W_{i,t}$ only includes the intercept. Model 2 is a Markov switching model for which only the autocorrelation covariates described in section 5.1 are included for the estimation of $\lambda$.
Models 3, 4 (and 5) include time-specific interest covariates and notably age. Age is encoded as time elapsed since the individual was admitted to ENA, on a range from 0 to 1 (with 1 corresponding to the entire time frame, 32 years). Age is modeled in a semi-parametric fashion, using B-splines of degree 2, with 5 degrees of freedom including an intercept, and boundary knots in 0 and 1, leading to g=1,,5g=1,...,5 covariates $S_{i,t}^{g}$ in the linear model (with the R package splines). We do this in order to saturate the model on the age dimension, so we can safely identify an effect related to cohort or period effects. Using splines works better than multiple powers of age: under this parameterization, parameters are less correlated, so the sampler mixes faster. As the spline basis already includes an intercept, no general intercept is included. Recall from our discussion in Section 5.1 that we have to use a more complex scheme to handle boundary effects for $\lambda$, and interact these five covariates with $\mathbb{I}_{A(1)=0}$ for all three models. More specifically, Model 3 adds a time-varying component depending on the age of individuals for $\gamma_{0}$ and $\gamma_{1}$. Model 4 adds a cohort effect for the estimation of $\gamma^{0}$, $\gamma^{1}$, and $\lambda$. Cohort is encoded as the time at which the individual was admitted to ENA, ranging between 0 and 1.
Finally, we examine the dependence between public-private paths and electoral cycles. We implement an additional Model 5, which is similar to Model 4, except for the fact that instead of a cohort effect, we estimate a period effect using a dummy variable $P_{i,t}$ with value $1$ if $t$ or $t+1$ is an election time point, and 0 else. Election time points correspond to Jan-Jun 1995, 2002, 2007, 2012, 2017, 2022 (French presidential elections). This dummy is shifted by one unit (so $t+1$ or $t+2$) for $\lambda$.
@@ 5.2.2 Analysis of Model Accuracy
We first analyse the output for models 1, 2, and 3. We examine both the posterior of parameters for simpler models and the quality of data augmentation by performing analyses where detailed data source information is left out for some observed individuals.
Parameters posterior. The simplicity of models 1 and 2 enables full reporting on the posterior of their coefficients, which we do in Table LABEL:tab:mod12 with their 0.05, 0.50 and 0.95 quantiles.
Table 3: Coefficients for Models 1 and 2.
$M_{1}$ $M_{1}$ $M_{1}$ $M_{2}$ $M_{2}$ $M_{2}$
@@
$\gamma_{1}$, (Intercept) -4.054 -3.938 -3.830 -3.667 -3.525 -3.391
$\lambda$, (Intercept) -1.510 -1.492 -1.474 -1.853 -1.786 -1.718
$\lambda$, $L$ -0.16 -0.134 -0.108
$\lambda$, $A(1)$ -0.133 0.088 0.315
@@
$\lambda$, $I(A(1)=0):L$ -0.173 -0.137 -0.101
Comparison between models 1 and 2 shows why there is a need for autoregressive coefficients and to handle boundary effects, as discussed in section 5.1. The main effect of adding autocorrelation coefficients from Model 1 to Model 2 can be observed in the difference between their median intercept for $\lambda$. For Model 1, the median intercept is at 1.492-1.492 ($I_{90}=[-1.51,-1.474]$), which essentially corresponds to an average probability of 0.22 for an individual to generate a trace at any point in time, which is higher than the one obtained for Model 2, at 1.786-1.786 ($I_{90}=[-1.853,-1.718]$), and which is more in line with what could be expected. All credible regions for autocorrelation coefficients except $A(1)$ do not contain zero despite that they are quite correlated by construction. Finally, the coefficient responsible for handling the boundary effect $\mathbb{I}_{A(1)=0}$ is very different from zero and negative, showing indeed that it is needed. Configurations without it also have very different values for all other coefficients.
Evaluation by reconstruction loss. We repeat the inference under models 1, 2, 3 but with some data withheld. In this experiment, we exclude from the data 20% of individuals in the detailed data source. We can thus assess the data augmentation quality could on points that are known but excluded from the analysis. We adopt a simple cross-entropy loss on the trajectory for such points $l_{i,t}=-(X_{i,t}^{\text{obs}}\log(X_{i,t}^{\text{pred}})+(1-X_{i,t}^{\text{obs}})\log(1-X_{i,t}^{\text{pred}})$), where predictions are computed as the average of observations generated in the Gibbs step over the last $300$ iterations of the chain, with probability clipping at 0.01 and 0.99, and observations are taken from the detailed data source. For Model 1, we also implement a version where no information about the traces was used at all, which we denote Model 1b.
The average cross-entropy loss obtained for the different models is 0.39 for Model 1, 0.29 for Model 1b, 0.22 for Model 2, and 0.21 for Model 3. A first very clear conclusion is that Model 1, the simple HMM without autoregressive components, is badly suited for the task, as it performs substantially worse than all other models. We observe that better reconstruction is obtained by including both data sources and autoregressive components (models 2 and 3). The change in accuracy from Model 2 to Model 3 is small in size, but our interest is rather to perform hypothesis testing when including the additional age coefficients. We might obtain better reconstruction by adopting more complex models, but this would come at the cost of model interpretability and hypothesis testing, which is our main focus.